2. If the coefficient of friction between vehicles and the road is 0.4 and the radius of curvature of




a circular curve of the road is 30 m. What is the maximum speed allowable if there is no




banking?

1
Expert's answer
2022-01-20T10:07:35-0500

"F_a=F_{fr},"

"\frac{mv^2}r=\mu mg,"

"v=\sqrt{\mu gr}=10.8~\frac ms."